### v1

# Barrow holographic dark energy with varying exponent

Abstract
We construct Barrow holographic dark energy with varying exponent. Such an energy-scale-dependent behavior is typical in quantum field theory and quantum gravity under renormalization group considerations, however in the present scenario it has an additional justification, since in realistic cases one expects that Barrow entropy quantum-gravitational effects to be stronger at early times and to smooth out and disappear at late times. We impose specific, redshift-dependent ansätze for the Barrow running exponent, such as the linear, CPL-like, exponential, and trigonometric ones, and we investigate their cosmological behavior. We show that we can recover the standard thermal history of the universe, with the sequence of matter and dark energy epochs, in which the transition from deceleration to acceleration happens at $z\approx 0.65$, in agreement with observations. In the most realistic case of hyperbolic tangent ansatz, in which we can easily bound Barrow exponent inside its theoretically determined bounds 0 and 1 for all redshifts, we see that the dark-energy equation-of-state parameter can be quintessence like, or experience the phantom-divide crossing, while in the future it can either tend to the cosmological constant value or start increasing again. All these features reveal that Barrow holographic dark energy with varying exponent is not only theoretically more justified than the standard, constant-exponent case, but it leads to richer cosmological behavior too.
1 Introduction
The progression from the matter era to the accelerated expansion phase in the late universe is now a well-established phenomenon. While the cosmological constant is the simplest explanation, challenges related to its quantum-field-theoretical calculation and the possibility of a dynamical nature have prompted two primary approaches to construct extended scenarios. The first retains general relativity as the underlying gravitational theory and introduces the concept of dark energy [1, 2, 3] as the acceleration source. The second involves constructing extended gravity theories which provide the required richer structure [4, 5, 6, 7].
An alternative explanation for the origin of dark energy emerges through the cosmological application of the holographic principle [8, 9, 10]. This framework, rooted in the thermodynamics of black holes, connects the ultraviolet cutoff of a quantum field theory with the largest distance of the theory, a prerequisite for its applicability at large distances [11]. In a region where entropy is proportional to volume, the total energy should not exceed the mass of a black hole with the same radius, and this saturation leads to the extraction of a holographically originated vacuum energy, and thus to a form of holographic dark energy with a dynamic nature [12, 13]. The cosmological implications of holographic dark energy are intriguing [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22] and align with observations [23, 24, 25, 26, 27, 28]. Additionally, one has the advantage that holographic scenarios are free from potential pathologies that may appear in modified gravity [7]. Hence, holographic dark energy has spurred extensive research, leading to various extensions [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58].
One large class of holographic dark energy extension is obtained by changing the underlying entropy relation. In particular, the base models use the standard Bekenstein-Hawking entropy. However, since there are many extended entropies, that arise through various considerations, such as Tsallis non-additive entropy [59], Barrow quantum-gravity corrected entropy [60], Kaniadakis relativistic entropy [61, 62], power-law corrected entropy [63, 64] etc, one can respectively obtain Tsallis holographic dark energy [65, 66], Barrow holographic dark energy [67, 68, 69], Kaniadakis holographic dark energy [70, 71], power-law holographic dark energy [72], etc.
All the above extended-entropy scenarios incorporate a constant parameter that quantifies the deviation form standard entropy. However, in principle one expects a dynamical, energy-scale-dependent behavior for this parameter, aligning with the typical case observed in quantum field theory and quantum gravity under renormalization group applications. In particular, all parameters and coupling constants demonstrate a running nature in tandem with the energy scale, and in a cosmological framework this would effectively generate a time-dependence. A first investigation of holographic dark energy models with extended entropy with varying exponent was performed in [73] and it was shown that the running behavior can lead to interesting physical implications.
From the aforementioned extended entropies, Barrow entropy has been shown to lead to a very interesting cosmological phenomenology. It arises from quantum-gravitational phenomena on the horizon structure, parametrized by the single exponent $\Delta$ [60], where $\Delta=0$ corresponds to standard entropy and $\Delta=1$ to maximal deviation. Hence, it is natural to consider that at early times these quantum phenomena are more intense and thus $\Delta$ is closer to 1, while as time passes they smooth out and $\Delta$ tends to its standard value 0. In summary, a scenario of Barrow entropy with varying exponent has even more justification than the usual energy-scale dependence of coupling constants.
In the present work we are interested in studying Barrow holographic dark energy with varying exponent. In particular, we impose various parametrizations for the time-dependence, which is equivalent to redshift-dependence, and we examine their effect on the cosmological evolution. The plan of the work is the following. In Section 2 we review standard Barrow holographic dark energy, and then in Section 3 we construct the extended scenario where the Barrow exponent is varying. In Section 4 we impose specific ansätze for the Barrow running exponent and we investigate their cosmological behavior. Finally, in Section 5 we summarize our results.
2 Standard Barrow holographic dark energy
In this section we briefly review the scenario of Barrow holographic dark energy following [67]. Barrow entropy is given by [60]
$S_{B}=\left(\frac{A}{A_{0}}\right)^{1+\Delta/2},$ (2.1)
where $A$ is the standard horizon area and $A_{0}$ the Planck area, while the exponent $\Delta$ quantifies the quantum-gravitational deformation. Incorporating (2.1) in the definition of the standard holographic dark energy, expressed as the inequality $\rho_{DE}L^{4}\leq S$, with $L$ the horizon length, and imposing that $S\propto A\propto L^{2}$ [13] will result in
$\rho_{DE}={C}L^{\Delta-2},$ (2.2)
with ${C}$ a parameter with dimensions $[L]^{-2-\Delta}$. In case where $\Delta=0$, expression (2.2) recovers standard holographic dark energy $\rho_{DE}=3c^{2}M_{p}^{2}L^{-2}$ (here $M_{p}$ is the Planck mass), where ${C}=3c^{2}M_{p}^{2}$ and $c^{2}$ the model parameter.
We focus on a flat homogeneous and isotropic Friedmann-Robertson-Walker (FRW) geometry with metric
$ds^{2}=-dt^{2}+a^{2}(t)\delta_{ij}dx^{i}dx^{j}\,,$ (2.3)
where $a(t)$ is the scale factor. Furthermore, we consider that the universe is filled with the matter perfect fluid, as well as with the Barrow holographic dark energy [67]
$\rho_{DE}={C}R_{h}^{\Delta-2},$ (2.4)
where the horizon length $L$ in (2.2) is substituted by the future event horizon $R_{h}$[12], given by
$R_{h}\equiv a\int_{t}^{\infty}\frac{dt}{a}=a\int_{a}^{\infty}\frac{da}{Ha^{2}},$ (2.5)
with $H\equiv\dot{a}/a$ the Hubble parameter. We then obtain the two Friedmann equations
$3M_{p}^{2}H^{2}$ $=$ $\ \rho_{m}+\rho_{DE}$ (2.6)
$-2M_{p}^{2}\dot{H}$ $=$ $\rho_{m}+p_{m}+\rho_{DE}+p_{DE},$ (2.7)
with $\rho_{m}$ and $p_{m}$ the energy density and pressure of matter and $p_{DE}$ the pressure of Barrow holographic dark energy.
The matter sector is conserved, namely it satisfies the continuity equation $\dot{\rho}_{m}+3H(\rho_{m}+p_{m})=0$. Considering matter to be dust, namely imposing $p_{m}=0$, leads to $\rho_{m}=\rho_{m0}/a^{3}$, with $\rho_{m0}$ the present matter energy density, i.e. at $a_{0}=1$ (in the following the subscript “0” denotes the value of the corresponding quantity at present). Finally, we focus on physically interesting observables such as the density parameters
$\Omega_{m}\equiv\frac{1}{3M_{p}^{2}H^{2}}\rho_{m}$ (2.8)
$\Omega_{DE}\equiv\frac{1}{3M_{p}^{2}H^{2}}\rho_{DE},$ (2.9)
as well as on the effective dark-energy equation-of-state parameter
$w_{DE}\equiv\frac{p_{DE}}{\rho_{DE}}.$ (2.10)
Combining the aforementioned density parameters with equations (2.4) and (2.5), leads to [67]
$\int_{x}^{\infty}\frac{dx}{Ha}=\frac{1}{a}\left(\frac{{C}}{3M_{p}^{2}H^{2}\Omega_{DE}}\right)^{\frac{1}{2-\Delta}},$ (2.11)
with $x\equiv\ln a$. Additionally, substituting $\rho_{m}=\rho_{m0}/a^{3}$ into (2.8), results to $\Omega_{m}=\Omega_{m0}H_{0}^{2}/(a^{3}H^{2})$, from which, using the Friedmann equation $\Omega_{m}+\Omega_{DE}=1$, we obtain
$\frac{1}{Ha}=\frac{\sqrt{a(1-\Omega_{DE})}}{H_{0}\sqrt{\Omega_{m0}}}.$ (2.12)
Inserting (2.12) into equation (2.11) we acquire the expression
$\int_{x}^{\infty}\frac{dx}{H_{0}\sqrt{\Omega_{m0}}}\sqrt{a(1-\Omega_{DE})}=\frac{1}{a}\left(\frac{{C}}{3M_{p}^{2}H^{2}\Omega_{DE}}\right)^{\frac{1}{2-\Delta}}.$ (2.13)
Taking the derivative of (2.13) with respect to $x=\ln a$ results to
$\!\!\!\!\!\!\!\!\!\!\!\!\!\frac{\Omega_{DE}^{\prime}}{\Omega_{DE}(1-\Omega_{DE})}=\Delta+1+Q(1-\Omega_{DE})^{\frac{\Delta}{2(\Delta-2)}}(\Omega_{DE})^{\frac{1}{2-\Delta}}e^{\frac{3\Delta}{2(\Delta-2)}x},$ (2.14)
with
$Q\equiv(2-\Delta)\left(\frac{{C}}{3M_{p}^{2}}\right)^{\frac{1}{\Delta-2}}\left(H_{0}\sqrt{\Omega_{m0}}\right)^{\frac{\Delta}{2-\Delta}},$ (2.15)
and where primes denote derivatives with respect to $x$. This is the differential equation determining the evolution of Barrow holographic dark energy, whose solution provides $\Omega_{DE}$. Finally, since from (2.4) we acquire $\dot{\rho}_{DE}=(\Delta-2){C}R_{h}^{\Delta-3}\dot{R}_{h}$, with $\dot{R}_{h}$ calculated using (2.5) as $\dot{R}_{h}=HR_{h}-1$, and since $\dot{\rho}_{DE}+3H(\rho_{DE}+p_{DE})=0$, using (2.14) we finally extract [67]
$w_{DE}=-\frac{1\!+\!\Delta}{3}-\frac{Q}{3}(\Omega_{DE})^{\frac{1}{2-\Delta}}(1\!-\!\Omega_{DE})^{\frac{\Delta}{2(\Delta-2)}}e^{\frac{3\Delta}{2(2-\Delta)}x}.$ (2.16)
All the above expressions, for $\Delta=0$ recover standard holographic dark energy with $w_{DE}|_{\Delta=0}=-\frac{1}{3}-\frac{2}{3}\sqrt{\frac{3M_{p}^{2}\Omega_{DE}}{C}}$ [12, 13]. The scenario of Barrow holographic dark energy has interesting cosmological implications, and has been studied in detail in the literature [67, 74, 75, 76, 77, 69, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90].
3 Barrow holographic dark energy with varying exponent
In this section we will investigate the cosmological scenario of Barrow holographic dark energy with varying exponent $\Delta$, namely we consider that $\Delta\equiv\Delta(x)$. In this case, the analysis of the previous section remains the same up to (2.13). However, taking its derivative requires to consider also the derivative of $\Delta$, which will yield $\Delta^{\prime}$ terms in the equations. In particular, differentiating (2.13) with respect to $x\equiv\ln a$, leads to
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\frac{\Omega_{DE}^{\prime}}{\Omega_{DE}\left(1-\Omega_{DE}\right)}=\sqrt{\Omega_{DE}}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\Delta}{2(\Delta-2)}}(2-\Delta)+\Delta+1$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\log{\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]}^{\frac{\Delta^{{}^{\prime}}}{\Delta-2}},$ (3.1)
with
$Q=Q(x)\equiv\frac{Ce^{3x}}{3M_{p}^{2}H_{0}^{2}\Omega_{m0}}.$ (3.2)
Equation (3) is the one that determines the evolution of Barrow holographic dark energy for dust matter in a flat universe, where primes denote derivatives with respect to $x$.
Additionally, let us calculate the equation-of-state parameter $w_{DE}$ for this general scenario. We differentiate (2.4), which leads to $\dot{\rho}_{DE}={C}R_{h}^{\Delta-2}\left[\log{R}_{h}\,\dot{\Delta}+\frac{\dot{R_{h}}(\Delta-2)}{R_{h}}\right]$, with $\dot{R}_{h}=HR_{h}-1$. Taking into account the dark energy conservation equation $\dot{\rho}_{DE}+3H\rho_{DE}(1+w_{DE})=0$ and the aforementioned relations, we result to
$(\Delta-2){C}\left(\frac{\rho_{DE}}{{C}}\right)^{\frac{\Delta-3}{\Delta-2}}\left[H\left(\frac{\rho_{DE}}{{C}}\right)^{\frac{1}{\Delta-2}}-1\right]$
$+\rho_{DE}\dot{\Delta}\log\left(\frac{\rho_{DE}}{C}\right)^{\frac{1}{\Delta-2}}+3H\rho_{DE}(1+w_{DE})=0.$ (3.3)
Hence, inserting $H$ from (2.12), and using (2.9) we finally obtain
$w_{DE}=-\frac{\Delta\!+\!1}{3}+\frac{(\Delta\!-\!2)\sqrt{\Omega_{DE}}}{3}\left(\!\frac{{C}}{3M_{p}^{2}}\!\right)^{\!-\frac{1}{2}}\!\left[\frac{Q(1\!-\!\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\Delta}{2(\Delta\!-\!2)}}-\frac{\Delta^{{}^{\prime}}}{3}\log{\left[\frac{Q(1\!-\!\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{1}{2\!-\!\Delta}}}.$ (3.4)
In the case where $\Delta=const.$ we recover equation (2.16) of the previous section.
4 Cosmological evolution
In the previous section we extracted the necessary cosmological equations, which can describe the evolution in the scenario at hand. In order to investigate in detail the cosmological behavior, in what follows it is more convenient to use the redshift $z$ instead of $x$ through the relation $x\equiv\ln a=-\ln(1+z)$. Hence, the derivative of a function $f$ in terms of $x$ easily becomes derivative in terms of $z$ through $f^{\prime}=-(1+z)\frac{df}{dz}$.
We will consider several ansätze for the Barrow running exponent, which will be given by the general expression $\Delta(z)=\alpha+\beta f(z)$, where $\alpha$, $\beta$ are constants and $f(z)$ a function of the redshift parameter $z$. Nevertheless, we impose that the exponent $\Delta$ acquires values away from the standard 00-value at early times, while as time passes it tends closer to the standard value $\Delta=0$.
Since the differential equation (3) cannot be solved analytically, we will elaborate it numerically. For initial conditions we impose $\Omega_{DE}(x=-\ln(1+z)=0)\equiv\Omega_{DE0}\approx 0.7$ and therefore $\Omega_{m}(x=-\ln(1+z)=0)\equiv\Omega_{m0}\approx 0.3$ in agreement with observations [91].
4.1 Linear case
As a first example, let us consider the linear ansatz $\Delta(z)=\alpha+\beta z$. Inserting it into equation (3) we acquire
$\!\!\!\!\!\!\!\!\!\!\!\!\!\frac{-(1+z)}{\Omega_{DE}\left(1-\Omega_{DE}\right)}\frac{d\Omega_{DE}}{dz}=\sqrt{\Omega_{DE}}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta z}{2(\alpha+\beta z-2)}}(2-\alpha-\beta z)$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \,\ \ \ \ \ \ \ \ \ \ \ \ +\log{\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]}^{\frac{-\beta(1+z)}{\alpha+\beta z-2}}+\alpha+\beta z+1,\hskip 28.00006pt$ (4.1)
and consequently equation (3.4) becomes
$\!\!\!\!\!\!\!\!\!\!\!\!\!w_{DE}=-\frac{\alpha+\beta z+1}{3}+\frac{(\alpha+\beta z-2)\sqrt{\Omega_{DE}}}{3}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta z}{2(\alpha+\beta z-2)}}$
$\ \ \ +\frac{(1+z)\beta}{3}\log{\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{1}{2-\alpha-\beta z}}}.$ (4.2)
Figure 1: Barrow holographic dark energy in the linear case where $\Delta(z)=\alpha+\beta z$. Upper graph: The dark energy density parameter $\Omega_{DE}$ (blue-solid) and the matter density parameter $\Omega_{m}$ (red-dashed), as a function of the redshift $z$, for $C=1$ and $\alpha=10^{-3}$, $\beta=10^{-4}$, in units where $M^{2}_{p}=1$. Middle graph: The corresponding dark energy equation-of-state parameter $w_{DE}$. Lower graph: The corresponding deceleration parameter $q$. In all graphs we have set $\Omega_{DE}(x=-\ln(1+z)=0)\equiv\Omega_{DE0}\approx 0.7$, in agreement with observations.
In the upper graph of Fig. 1, we depict the evolution of the dark energy and matter density parameters in terms of redshift. As we can see, we acquire the usual thermal history of the universe, with the sequence of matter and dark energy epochs, and in the far future, namely at $z\rightarrow-1$, the universe results asymptotically to a complete dark energy dominated phase. Furthermore, in the middle graph we depict the evolution of the dark energy equation-of-state parameter, which is determined by (4.1). As we observe, the value of $w_{DE}$ at present is around $-1$ in agreement with observations, lying in the quintessence regime, while in the future it asymptotically goes to the cosmological constant value. Lastly, in the lower graph we depict the deceleration parameter
$q\equiv-1-\frac{\dot{H}}{H^{2}}=\frac{1}{2}+\frac{3}{2}w_{DE}\Omega_{DE}.$ (4.3)
We can see the transition from deceleration to acceleration at $z\approx 0.65$, in agreement with observational data. Note that, as required, the parameter $\beta$ should be suitably smaller than $\alpha$ in order for $\Delta$ to remain between 0 and 1. However, since there will be always a $z$ in which this will not be the case, it is clear that this ansatz is valid only at small redshifts. Hence, in the following we proceed to the investigation of ansätze with full applicability.
4.2 CPL-like case
Inspired by the Chevallier-Polarski-Linder (CPL) parameterization for the dark-energy equation of state [92, 93], we consider a varying exponent of the form $\Delta(z)=\alpha+\frac{\beta z}{z+1}$, which can always be less than 1 for arbitrarily large redshifts. In this case substituting into (3) we obtain
$\!\!\!\!\!\!\!\!\!\frac{-(1+z)}{\Omega_{DE}\left(1-\Omega_{DE}\right)}\frac{d\Omega_{DE}}{dz}=\sqrt{\Omega_{DE}}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha(z+1)+\beta z}{2(\alpha-2)(z+1)+2\beta z}}\left[2\!-\!\alpha\!-\!\frac{\beta z}{z+1}\right]$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \,\ \ \ \ \ \ \ +\log{\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]}^{-\frac{\beta}{(\alpha-2)(z+1)+\beta z}}+\alpha+\frac{\beta z}{z+1}+1,\hskip 28.00006pt$ (4.4)
while equation (3.4) gives
$\!\!\!\!\!\!\!\!\!\!w_{DE}=-\frac{\alpha+\frac{\beta z}{z+1}+1}{3}+\frac{\left(\alpha+\frac{\beta z}{z+1}-2\right)\sqrt{\Omega_{DE}}}{3}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha(z+1)+\beta z}{2(\alpha-2)(z+1)+2\beta z}}$
$\ \ \ \ \ -\frac{\beta}{3(z+1)}\log{\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{z+1}{(2-\alpha)(z+1)-\beta z}}}.$ (4.5)
The behavior of dark energy and matter density parameters, of the deceleration parameter, and of the dark-energy equation-of-state parameter, is similar to the one of the linear case. Nevertheless, although very efficient in describing the past Universe evolution, the CPL-like parametrization is not suitable for the future evolution, since the Barrow exponent will become larger than 1 or smaller than 0, and hence in the following we investigate more realistic cases.
4.3 Exponential case
One ansatz that can be suitable for all redshifts is the exponential one, namely $\Delta(z)=\alpha+\beta e^{-\lambda z}$, where $\lambda$ is a constant. In this case substituting into (3) we obtain
$\!\!\!\!\!\!\!\!\!\!\!\!\!\frac{-(1+z)}{\Omega_{DE}\left(1-\Omega_{DE}\right)}\frac{d\Omega_{DE}}{dz}=\sqrt{\Omega_{DE}}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta e^{-\lambda z}}{2(\alpha+\beta e^{-\lambda z}-2)}}[2-(\alpha+\beta e^{-\lambda z})]$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \,\ \ \ \ \ +\log{\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]}^{\frac{(1+z)\beta\lambda e^{-\lambda z}}{\alpha+\beta e^{-\lambda z}-2}}+\alpha+\beta e^{-\lambda z}+1,\hskip 28.00006pt$ (4.6)
while equation (3.4) gives
$\!\!\!\!\!\!\!\!\!\!w_{DE}=-\frac{\alpha+\beta e^{-\lambda z}+1}{3}+\frac{(\alpha+\beta e^{-\lambda z}-2)\sqrt{\Omega_{DE}}}{3}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta e^{-\lambda z}}{2(\alpha+\beta e^{-\lambda z}-2)}}$
$\ \ \ \ \ -\frac{(1+z)\beta\lambda e^{-\lambda z}}{3}\log{\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{1}{2-\alpha-\beta e^{-\lambda z}}}}.$ (4.7)
Figure 2: Barrow holographic dark energy in the exponential case with $\Delta(z)=\alpha+\beta e^{-\lambda z}$. Upper graph: The dark energy density parameter $\Omega_{DE}$ (blue-solid) and the matter density parameter $\Omega_{m}$ (red-dashed), as a function of the redshift $z$, for $C=1$ and $\alpha=10^{-3}$, $\beta=10^{-4}$, $\lambda=5\cdot 10^{-5}$, in units where $M^{2}_{p}=1$. Middle graph: The corresponding dark energy equation-of-state parameter $w_{DE}$. Lower graph: The corresponding deceleration parameter $q$. In all graphs we have set $\Omega_{DE}(x=-\ln(1+z)=0)\equiv\Omega_{DE0}\approx 0.7$, in agreement with observations.
Fig. 2 shows the evolution of the dark energy and matter density parameters, and similarly to the linear case we acquire the usual thermal history of the universe. Additionally, from the middle graph we observe that $w_{DE}$ lies in the quintessence regime and its present value is around $-1$ according to observations. Finally, in the lower graph we depict the deceleration parameter $q$, where the transition from deceleration to acceleration in this case happens at $z\approx 0.65$.
4.4 Hyperbolic Tangent case
One ansatz that can be very suitable for the purpose of Barrow holographic dark energy with varying exponent is the hyperbolic tangent one, since in this case we can immediately bound $\Delta$ between 0 and 1 for all redshifts, obtaining easily a case where $\Delta(z)$ is 1 at early times while it becomes 0 at intermediate and late times, as well as in the future.
Figure 3: Barrow holographic dark energy in the Hyperbolic Tangent case with $\Delta(z)=\alpha+\beta\tanh({\gamma z})$, for $\alpha=\beta=\frac{1}{2}$. Upper graph: The dark energy density parameter $\Omega_{DE}$ (blue-solid) and the matter density parameter $\Omega_{m}$ (red-dashed), as a function of the redshift $z$, for $C=1$ and $\gamma=0.001$, in units where $M^{2}_{p}=1$. Middle graph: The corresponding dark energy equation-of-state parameter $w_{DE}$. Lower graph: The corresponding deceleration parameter $q$. In all graphs we have set $\Omega_{DE}(x=-\ln(1+z)=0)\equiv\Omega_{DE0}\approx 0.7$, in agreement with observations.
We consider the form $\Delta(z)=\alpha+\beta\tanh({\gamma z})$. Inserting into equation (3) yields
$\!\!\!\!\frac{-(1+z)}{\Omega_{DE}\left(1-\Omega_{DE}\right)}\frac{d\Omega_{DE}}{dz}=\sqrt{\Omega_{DE}}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta\tanh({\gamma z})}{2[\alpha+\beta\tanh({\gamma z})-2)]}}[2-\alpha-\beta\tanh({\gamma z})]$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\log{\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]}^{\frac{-(1+z)\beta\gamma\text{sech}^{2}({\gamma z})}{\alpha+\beta\tanh({\gamma z})-2}}+\alpha+\beta\tanh({\gamma z})+1,\hskip 28.00006pt$ (4.8)
and
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!w_{DE}=\frac{[\alpha\!+\!\beta\tanh({\gamma z})\!-\!2]\sqrt{\Omega_{DE}}}{3}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\!\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta\tanh({\gamma z})}{2[\alpha\!+\!\beta\tanh({\gamma z})-2)]}}$
$\ \ \ \,-\frac{\alpha\!+\!\beta\tanh({\gamma z})\!+\!1}{3}+\frac{(1+z)\beta\gamma\,\text{sech}^{2}({\gamma z})}{3}\log{\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{1}{2\!-\!\alpha\!-\!\beta\tanh{\gamma z}}}}.$ (4.9)
In the following we set $\alpha=\beta=\frac{1}{2}$ which fixes the largest value of $\Delta(z)$ to 1 and the smallest to 0, in consistency with Barrow entropy. Hence the only model parameter is $\gamma$, which determines how fast the transition from 1 to 0 takes place.
Figure 4: The evolution of the dark-energy equation-of-state parameter for Barrow holographic dark energy in the Hyperbolic Tangent case with $\Delta(z)=\alpha+\beta\tanh({\gamma z})$, for $\alpha=\beta=\frac{1}{2}$. The various curves correspond to $\gamma=0.01$ (black solid), $\gamma=0.05$ (green dashed), $\gamma=0.1$ (red dotted), $\gamma=0.2$ (orange dashed-dotted), $\gamma=0.3$ (magenta dashed-dotted-dotted), $\gamma=0.4$ (cyan short-dashed), and $\gamma=0.5$ (blue short dashed-dotted), in units where $M^{2}_{p}=1$. In all cases we have imposed $\Omega_{DE}(x=-\ln(1+z)=0)\equiv\Omega_{DE0}\approx 0.7$ at present, in agreement with observations.
Figure 4: The evolution of the dark-energy equation-of-state parameter for Barrow holographic dark energy in the Hyperbolic Tangent case with $\Delta(z)=\alpha+\beta\tanh({\gamma z})$, for $\alpha=\beta=\frac{1}{2}$. The various curves correspond to $\gamma=0.01$ (black solid), $\gamma=0.05$ (green dashed), $\gamma=0.1$ (red dotted), $\gamma=0.2$ (orange dashed-dotted), $\gamma=0.3$ (magenta dashed-dotted-dotted), $\gamma=0.4$ (cyan short-dashed), and $\gamma=0.5$ (blue short dashed-dotted), in units where $M^{2}_{p}=1$. In all cases we have imposed $\Omega_{DE}(x=-\ln(1+z)=0)\equiv\Omega_{DE0}\approx 0.7$ at present, in agreement with observations.
In the upper graph of Fig. 3 we depict the behavior of the dark energy and matter density parameters. As we can see, we acquire the sequence of matter and dark energy epochs, and in the far future the universe results asymptotically to a complete dark-energy dominated phase. Furthermore, the dark energy equation-of-state parameter, in this specific example lies in the quintessence regime, while at present times and in the future it is around $-1$. Lastly, the deceleration parameter $q$ experiences the transition from deceleration to acceleration at around $z\approx 0.65$.
Since the hyperbolic tangent form is the most realistic case, we explore it further. In particular, we are interested in examining what is the effect of the single parameter $\gamma$ on the dark-energy equation-of-state parameter. In Fig. 4, we depict the evolution of $w_{DE}$, for various values of the $\gamma$. A general observation is that for increasing $\gamma$ the value of $w_{DE}$ decreases. Additionally, one can see that for some cases it can experience the phantom-divide crossing. Interestingly enough, in the future $w_{DE}$ can either tend to $-1$ or even start increasing again. These features reveals the capabilities of the model, and the advantages of Barrow holographic dark energy with running exponent comparing to standard Barrow holographic dark energy.
4.5 Trigonometric functions case
Finally, inspired by the oscillatory parametrizations of the dark-energy equation-of-state parameters [94], one could think of richer cases, where the Barrow exponent oscillates around a mean positive value close to 0. One such case could be the $\Delta(z)=\alpha+\beta\sin{z}$ with $0<\beta<\alpha$, which leads to
$\!\!\!\!\!\!\!\!\!\!\!\!\!\frac{-(1+z)\Omega_{DE}^{\prime}}{\Omega_{DE}\left(1-\Omega_{DE}\right)}=\sqrt{\Omega_{DE}}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta\sin{z}}{2(\alpha+\beta\sin{z}-2)}}(2-\alpha-\beta\sin{z})$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \,\ +\log{\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]}^{\frac{-(1+z)\beta\cos{z}}{\alpha+\beta\sin{z}-2}}+\alpha+\beta\sin{z}+1,\hskip 32.00002pt$ (4.10)
$\!\!\!\!\!\!\!\!w_{DE}=-\frac{\alpha+\beta\sin{z}+1}{3}+\frac{(\alpha+\beta\sin{z}-2)\sqrt{\Omega_{DE}}}{3}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta\sin{z}}{2(\alpha+\beta\sin{z}-2)}}$
$\ \ \ \ \ \ +\frac{(1+z)\beta\cos{z}}{3}\log{\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{1}{2-\alpha-\beta\sin{z}}}}.$ (4.11)
Alternatively, one could consider $\Delta(z)=\alpha+\beta\cos{z}$, obtaining
$\!\!\!\!\!\!\!\!\!\!\!\!\!\frac{-(1+z)\Omega_{DE}^{\prime}}{\Omega_{DE}\left(1-\Omega_{DE}\right)}=\sqrt{\Omega_{DE}}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta\cos{z}}{2(\alpha+\beta\cos{z}-2)}}(2-\alpha-\beta\cos{z})$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\log{\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]}^{\frac{(1+z)\beta\sin{z}}{\alpha+\beta\cos{z}-2}}+\alpha+\beta\cos{z}+1,\hskip 28.00006pt$ (4.12)
$\!\!\!\!\!\!\!w_{DE}=-\frac{\alpha+\beta\cos{z}+1}{3}+\frac{(\alpha+\beta\cos{z}-2)\sqrt{\Omega_{DE}}}{3}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta\cos{z}}{2(\alpha+\beta\cos{z}-2)}}$
$\ \ \ \ \ \ \,-\frac{(1+z)\beta\sin{z}}{3}\log{\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{1}{2-\alpha-\beta\cos{z}}}}$ (4.13)
The behavior of dark energy and matter density parameters, of the deceleration parameter, and of the dark-energy equation-of-state parameter, is similar to the one of the previous models.
5 Conclusions
In this work we constructed Barrow holographic dark energy with varying exponent. This scenario is an extension of holographic dark energy with Barrow entropy, where the involved Barrow exponent, which quantifies the deviation from standard Bekenstein-Hawking entropy, is allowed to present a running behavior. Such an energy-scale-dependent behavior is typical in quantum field theory and quantum gravity under renormalization group considerations, however in the present scenario it has an additional justification, if not necessity, since in realistic cases one expects that Barrow entropy quantum-gravitational effects to be stronger at early times and to smooth out and disappear at late times. Hence, in the cosmological framework of an expanding Universe, this would effectively generate Barrow-exponent with time-dependence.
After constructing the extended scenario we imposed specific ansätze for the Barrow running exponent and we investigated their cosmological behavior. We started with the linear case, where we showed that we can recover the standard thermal history of the universe, with the sequence of matter and dark energy epochs. Additionally, we saw that the dark-energy equation-of-state parameter $w_{DE}$ lies in the quintessence regime and tends to $-1$ in the future, where the universe results asymptotically to a complete dark energy dominated phase. Finally, from the behavior of the deceleration parameter we saw that the transition from deceleration to acceleration happens at $z\approx 0.65$, in agreement with observations.
The linear ansatz is suitable only for small redshifts, since there will be always a redshift in which the Barrow exponent could exceed the theoretically determined bounds. The CPL-like ansatz is applicable at all redshifts in the past, however is still not suitable for the future evolution. Hence, we proceeded to the examination of a more realistic case, namely the exponential ansatz. In this model, we obtained a similar behavior, with a sequence of matter and dark-energy epochs and the transition to acceleration at $z\approx 0.65$ too.
Nevertheless, the most suitable ansatz for the purpose of Barrow holographic dark energy with varying exponent is the hyperbolic tangent one, since in this case we can easily bound Barrow exponent between 0 and 1 for all redshifts, obtaining a scenario where $\Delta(z)$ is 1 at early times while it becomes 0 at intermediate and late times, as well as in the future. This scenario can also recover the thermal history of the Universe, and the recent transition to acceleration. Furthermore, we investigated the effect of the single parameter $\gamma$ on the dark-energy equation-of-state parameter. We showed that for increasing $\gamma$ the value of $w_{DE}$ decreases, while for some cases it can experience the phantom-divide crossing, and that in the future $w_{DE}$ can either tend to $-1$ or start increasing again. Lastly, for completeness, we also presented the case where the running Barrow exponent has a trigonometric form.
All these features reveal that Barrow holographic dark energy with varying exponent is not only theoretically more justified than the standard, constant-exponent case, but it leads to richer cosmological behavior too. It would be interesting to perform observational confrontation with data from Supernovae type Ia (SNIa), Baryonic Acoustic Oscillations (BAO), Cosmic Microwave Background (CMB), and Cosmic Chronometers (CC) probes, in order to extract constraints on the involved parameters, and apply information criteria in order to compare the scenario with the concordance $\Lambda$CDM cosmology. Such a full observational analysis is left for a future project.
Acknowledgments
M.P. is supported by the Basic Research program of the National Technical University of Athens (NTUA, PEVE) 65232600-ACT-MTG: Alleviating Cosmological Tensions Through Modified Theories of Gravity. The authors acknowledge the contribution of the LISA CosWG and of the COST Action CA21136 “Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse)”.

### v2

@@
 Additionally, one has the advantage that holographic scenarios are free from potential pathologies that may appear in modified gravity [7].
-Hence, holographic dark energy has spurred extensive research, leading to various extensions [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58].
+Hence, holographic dark energy has spurred extensive research, leading to various extensions [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60].
 One large class of holographic dark energy extension is obtained by changing the underlying entropy relation.
@@
 In particular, all parameters and coupling constants demonstrate a running nature in tandem with the energy scale, and in a cosmological framework this would effectively generate a time-dependence.
-A first investigation of holographic dark energy models with extended entropy with varying exponent was performed in [73] and it was shown that the running behavior can lead to interesting physical implications.
-From the aforementioned extended entropies, Barrow entropy has been shown to lead to a very interesting cosmological phenomenology.
-It arises from quantum-gravitational phenomena on the horizon structure, parametrized by the single exponent $\Delta$ [60], where $\Delta=0$ corresponds to standard entropy and $\Delta=1$ to maximal deviation.
-Hence, it is natural to consider that at early times these quantum phenomena are more intense and thus $\Delta$ is closer to 1, while as time passes they smooth out and $\Delta$ tends to its standard value 0.
+A first investigation of Tsallis holographic dark energy models with extended entropy with varying exponent was performed in [75] and it was shown that the running behavior can lead to interesting physical implications.
+Nevertheless, in the case of Barrow entropy one has an additional necessity for the varying behavior.
+Since it arises from quantum-gravitational phenomena on the horizon structure, parametrized by the single exponent $\Delta$ [62], where $\Delta=0$ corresponds to standard entropy and $\Delta=1$ to maximal deviation, it is hard to justify why in the recent Universe one has significant quantum-gravitational phenomena in its horizon that could lead to significant deviation from standard, i.e.
+classical entropy.
+Hence, it is more natural to consider that at early times these quantum phenomena are more intense and thus $\Delta$ is closer to 1, while as time passes they smooth out and $\Delta$ tends to its standard value 0.
+We mention here that such a scenario is still Barrow entropy, since $\Delta$ does arise from quantum gravitational phenomena, however one goes beyond the first approximation of [62] in which these phenomena are constant throughout the Universe evolution, and examines more realistic scenarios in which quantum gravitational phenomena are stronger at early times and smooth out as the system becomes larger.
 In summary, a scenario of Barrow entropy with varying exponent has even more justification than the usual energy-scale dependence of coupling constants.
-In the present work we are interested in studying Barrow holographic dark energy with varying exponent.
+Since Barrow entropy has been shown to lead to a very interesting cosmological phenomenology, in the present work we are interested in studying Barrow holographic dark energy with varying exponent.
 In particular, we impose various parametrizations for the time-dependence, which is equivalent to redshift-dependence, and we examine their effect on the cosmological evolution.
@@
 In particular, differentiating (2.13) with respect to $x\equiv\ln a$, leads to
-$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\frac{\Omega_{DE}^{\prime}}{\Omega_{DE}\left(1-\Omega_{DE}\right)}=\sqrt{\Omega_{DE}}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\Delta}{2(\Delta-2)}}(2-\Delta)+\Delta+1$
-$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\log{\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]}^{\frac{\Delta^{{}^{\prime}}}{\Delta-2}},$ (3.1)
+$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\frac{\Omega_{DE}^{\prime}}{\Omega_{DE}\left(1-\Omega_{DE}\right)}=\sqrt{\Omega_{DE}}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\Delta}{2(\Delta-2)}}(2-\Delta)+\Delta+1$
+$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\log{\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]}^{\frac{\Delta^{{}^{\prime}}}{\Delta-2}},$ (3.1)
 with
-$Q=Q(x)\equiv\frac{Ce^{3x}}{3M_{p}^{2}H_{0}^{2}\Omega_{m0}}.$ (3.2)
+$P=P(x)\equiv\frac{Ce^{3x}}{3M_{p}^{2}H_{0}^{2}\Omega_{m0}}.$ (3.2)
 Equation (3) is the one that determines the evolution of Barrow holographic dark energy for dust matter in a flat universe, where primes denote derivatives with respect to $x$.
@@
 Hence, inserting $H$ from (2.12), and using (2.9) we finally obtain
-$w_{DE}=-\frac{\Delta\!+\!1}{3}+\frac{(\Delta\!-\!2)\sqrt{\Omega_{DE}}}{3}\left(\!\frac{{C}}{3M_{p}^{2}}\!\right)^{\!-\frac{1}{2}}\!\left[\frac{Q(1\!-\!\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\Delta}{2(\Delta\!-\!2)}}-\frac{\Delta^{{}^{\prime}}}{3}\log{\left[\frac{Q(1\!-\!\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{1}{2\!-\!\Delta}}}.$ (3.4)
+$w_{DE}=-\frac{\Delta\!+\!1}{3}+\frac{(\Delta\!-\!2)\sqrt{\Omega_{DE}}}{3}\left(\!\frac{{C}}{3M_{p}^{2}}\!\right)^{\!-\frac{1}{2}}\!\left[\frac{P(1\!-\!\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\Delta}{2(\Delta\!-\!2)}}-\frac{\Delta^{{}^{\prime}}}{3}\log{\left[\frac{P(1\!-\!\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{1}{2\!-\!\Delta}}}.$ (3.4)
 In the case where $\Delta=const.$ we recover equation (2.16) of the previous section.
@@
 Nevertheless, we impose that the exponent $\Delta$ acquires values away from the standard 00-value at early times, while as time passes it tends closer to the standard value $\Delta=0$.
+We mention here that we parameterize the way that quantum gravitational phenomena smooth out, i.e $\Delta$ as a function of redshift, in a purely phenomenological manner, through various functions.
+Definitely each of these functions at the fundamental level should correspond to a particular way that quantum gravitational phenomena behave.
 Since the differential equation (3) cannot be solved analytically, we will elaborate it numerically.
@@
 As a first example, let us consider the linear ansatz $\Delta(z)=\alpha+\beta z$. Inserting it into equation (3) we acquire
-$\!\!\!\!\!\!\!\!\!\!\!\!\!\frac{-(1+z)}{\Omega_{DE}\left(1-\Omega_{DE}\right)}\frac{d\Omega_{DE}}{dz}=\sqrt{\Omega_{DE}}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta z}{2(\alpha+\beta z-2)}}(2-\alpha-\beta z)$
-$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \,\ \ \ \ \ \ \ \ \ \ \ \ +\log{\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]}^{\frac{-\beta(1+z)}{\alpha+\beta z-2}}+\alpha+\beta z+1,\hskip 28.00006pt$ (4.1)
+$\!\!\!\!\!\!\!\!\!\!\!\!\!\frac{-(1+z)}{\Omega_{DE}\left(1-\Omega_{DE}\right)}\frac{d\Omega_{DE}}{dz}=\sqrt{\Omega_{DE}}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta z}{2(\alpha+\beta z-2)}}(2-\alpha-\beta z)$
+$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \,\ \ \ \ \ \ \ \ \ \ \ \ +\log{\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]}^{\frac{-\beta(1+z)}{\alpha+\beta z-2}}+\alpha+\beta z+1,\hskip 28.00006pt$ (4.1)
 and consequently equation (3.4) becomes
-$\!\!\!\!\!\!\!\!\!\!\!\!\!w_{DE}=-\frac{\alpha+\beta z+1}{3}+\frac{(\alpha+\beta z-2)\sqrt{\Omega_{DE}}}{3}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta z}{2(\alpha+\beta z-2)}}$
-$\ \ \ +\frac{(1+z)\beta}{3}\log{\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{1}{2-\alpha-\beta z}}}.$ (4.2)
+$\!\!\!\!\!\!\!\!\!\!\!\!\!w_{DE}=-\frac{\alpha+\beta z+1}{3}+\frac{(\alpha+\beta z-2)\sqrt{\Omega_{DE}}}{3}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta z}{2(\alpha+\beta z-2)}}$
+$\ \ \ +\frac{(1+z)\beta}{3}\log{\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{1}{2-\alpha-\beta z}}}.$ (4.2)
 Figure 1: Barrow holographic dark energy in the linear case where $\Delta(z)=\alpha+\beta z$.
@@
 In this case substituting into (3) we obtain
-$\!\!\!\!\!\!\!\!\!\frac{-(1+z)}{\Omega_{DE}\left(1-\Omega_{DE}\right)}\frac{d\Omega_{DE}}{dz}=\sqrt{\Omega_{DE}}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha(z+1)+\beta z}{2(\alpha-2)(z+1)+2\beta z}}\left[2\!-\!\alpha\!-\!\frac{\beta z}{z+1}\right]$
-$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \,\ \ \ \ \ \ \ +\log{\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]}^{-\frac{\beta}{(\alpha-2)(z+1)+\beta z}}+\alpha+\frac{\beta z}{z+1}+1,\hskip 28.00006pt$ (4.4)
+$\!\!\!\!\!\!\!\!\!\frac{-(1+z)}{\Omega_{DE}\left(1-\Omega_{DE}\right)}\frac{d\Omega_{DE}}{dz}=\sqrt{\Omega_{DE}}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha(z+1)+\beta z}{2(\alpha-2)(z+1)+2\beta z}}\left[2\!-\!\alpha\!-\!\frac{\beta z}{z+1}\right]$
+$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \,\ \ \ \ \ \ \ +\log{\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]}^{-\frac{\beta}{(\alpha-2)(z+1)+\beta z}}+\alpha+\frac{\beta z}{z+1}+1,\hskip 28.00006pt$ (4.4)
 while equation (3.4) gives
-$\!\!\!\!\!\!\!\!\!\!w_{DE}=-\frac{\alpha+\frac{\beta z}{z+1}+1}{3}+\frac{\left(\alpha+\frac{\beta z}{z+1}-2\right)\sqrt{\Omega_{DE}}}{3}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha(z+1)+\beta z}{2(\alpha-2)(z+1)+2\beta z}}$
-$\ \ \ \ \ -\frac{\beta}{3(z+1)}\log{\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{z+1}{(2-\alpha)(z+1)-\beta z}}}.$ (4.5)
+$\!\!\!\!\!\!\!\!\!\!w_{DE}=-\frac{\alpha+\frac{\beta z}{z+1}+1}{3}+\frac{\left(\alpha+\frac{\beta z}{z+1}-2\right)\sqrt{\Omega_{DE}}}{3}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha(z+1)+\beta z}{2(\alpha-2)(z+1)+2\beta z}}$
+$\ \ \ \ \ -\frac{\beta}{3(z+1)}\log{\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{z+1}{(2-\alpha)(z+1)-\beta z}}}.$ (4.5)
 The behavior of dark energy and matter density parameters, of the deceleration parameter, and of the dark-energy equation-of-state parameter, is similar to the one of the linear case.
@@
 One ansatz that can be suitable for all redshifts is the exponential one, namely $\Delta(z)=\alpha+\beta e^{-\lambda z}$, where $\lambda$ is a constant. In this case substituting into (3) we obtain
-$\!\!\!\!\!\!\!\!\!\!\!\!\!\frac{-(1+z)}{\Omega_{DE}\left(1-\Omega_{DE}\right)}\frac{d\Omega_{DE}}{dz}=\sqrt{\Omega_{DE}}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta e^{-\lambda z}}{2(\alpha+\beta e^{-\lambda z}-2)}}[2-(\alpha+\beta e^{-\lambda z})]$
-$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \,\ \ \ \ \ +\log{\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]}^{\frac{(1+z)\beta\lambda e^{-\lambda z}}{\alpha+\beta e^{-\lambda z}-2}}+\alpha+\beta e^{-\lambda z}+1,\hskip 28.00006pt$ (4.6)
+$\!\!\!\!\!\!\!\!\!\!\!\!\!\frac{-(1+z)}{\Omega_{DE}\left(1-\Omega_{DE}\right)}\frac{d\Omega_{DE}}{dz}=\sqrt{\Omega_{DE}}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta e^{-\lambda z}}{2(\alpha+\beta e^{-\lambda z}-2)}}[2-(\alpha+\beta e^{-\lambda z})]$
+$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \,\ \ \ \ \ +\log{\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]}^{\frac{(1+z)\beta\lambda e^{-\lambda z}}{\alpha+\beta e^{-\lambda z}-2}}+\alpha+\beta e^{-\lambda z}+1,\hskip 28.00006pt$ (4.6)
 while equation (3.4) gives
-$\!\!\!\!\!\!\!\!\!\!w_{DE}=-\frac{\alpha+\beta e^{-\lambda z}+1}{3}+\frac{(\alpha+\beta e^{-\lambda z}-2)\sqrt{\Omega_{DE}}}{3}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta e^{-\lambda z}}{2(\alpha+\beta e^{-\lambda z}-2)}}$
-$\ \ \ \ \ -\frac{(1+z)\beta\lambda e^{-\lambda z}}{3}\log{\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{1}{2-\alpha-\beta e^{-\lambda z}}}}.$ (4.7)
+$\!\!\!\!\!\!\!\!\!\!w_{DE}=-\frac{\alpha+\beta e^{-\lambda z}+1}{3}+\frac{(\alpha+\beta e^{-\lambda z}-2)\sqrt{\Omega_{DE}}}{3}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta e^{-\lambda z}}{2(\alpha+\beta e^{-\lambda z}-2)}}$
+$\ \ \ \ \ -\frac{(1+z)\beta\lambda e^{-\lambda z}}{3}\log{\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{1}{2-\alpha-\beta e^{-\lambda z}}}}.$ (4.7)
 Figure 2: Barrow holographic dark energy in the exponential case with $\Delta(z)=\alpha+\beta e^{-\lambda z}$.
@@
 We consider the form $\Delta(z)=\alpha+\beta\tanh({\gamma z})$. Inserting into equation (3) yields
-$\!\!\!\!\frac{-(1+z)}{\Omega_{DE}\left(1-\Omega_{DE}\right)}\frac{d\Omega_{DE}}{dz}=\sqrt{\Omega_{DE}}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta\tanh({\gamma z})}{2[\alpha+\beta\tanh({\gamma z})-2)]}}[2-\alpha-\beta\tanh({\gamma z})]$
-$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\log{\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]}^{\frac{-(1+z)\beta\gamma\text{sech}^{2}({\gamma z})}{\alpha+\beta\tanh({\gamma z})-2}}+\alpha+\beta\tanh({\gamma z})+1,\hskip 28.00006pt$ (4.8)
+$\!\!\!\!\frac{-(1+z)}{\Omega_{DE}\left(1-\Omega_{DE}\right)}\frac{d\Omega_{DE}}{dz}=\sqrt{\Omega_{DE}}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta\tanh({\gamma z})}{2[\alpha+\beta\tanh({\gamma z})-2)]}}[2-\alpha-\beta\tanh({\gamma z})]$
+$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\log{\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]}^{\frac{-(1+z)\beta\gamma\text{sech}^{2}({\gamma z})}{\alpha+\beta\tanh({\gamma z})-2}}+\alpha+\beta\tanh({\gamma z})+1,\hskip 28.00006pt$ (4.8)
 and
-$\!\!\!\!\!\!\!\!\!\!\!\!\!\!w_{DE}=\frac{[\alpha\!+\!\beta\tanh({\gamma z})\!-\!2]\sqrt{\Omega_{DE}}}{3}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\!\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta\tanh({\gamma z})}{2[\alpha\!+\!\beta\tanh({\gamma z})-2)]}}$
-$\ \ \ \,-\frac{\alpha\!+\!\beta\tanh({\gamma z})\!+\!1}{3}+\frac{(1+z)\beta\gamma\,\text{sech}^{2}({\gamma z})}{3}\log{\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{1}{2\!-\!\alpha\!-\!\beta\tanh{\gamma z}}}}.$ (4.9)
+$\!\!\!\!\!\!\!\!\!\!\!\!\!\!w_{DE}=\frac{[\alpha\!+\!\beta\tanh({\gamma z})\!-\!2]\sqrt{\Omega_{DE}}}{3}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\!\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta\tanh({\gamma z})}{2[\alpha\!+\!\beta\tanh({\gamma z})-2)]}}$
+$\ \ \ \,-\frac{\alpha\!+\!\beta\tanh({\gamma z})\!+\!1}{3}+\frac{(1+z)\beta\gamma\,\text{sech}^{2}({\gamma z})}{3}\log{\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{1}{2\!-\!\alpha\!-\!\beta\tanh{\gamma z}}}}.$ (4.9)
 In the following we set $\alpha=\beta=\frac{1}{2}$ which fixes the largest value of $\Delta(z)$ to 1 and the smallest to 0, in consistency with Barrow entropy.
 Hence the only model parameter is $\gamma$, which determines how fast the transition from 1 to 0 takes place.
-Figure 4: The evolution of the dark-energy equation-of-state parameter for Barrow holographic dark energy in the Hyperbolic Tangent case with $\Delta(z)=\alpha+\beta\tanh({\gamma z})$, for $\alpha=\beta=\frac{1}{2}$.
-The various curves correspond to $\gamma=0.01$ (black solid), $\gamma=0.05$ (green dashed), $\gamma=0.1$ (red dotted), $\gamma=0.2$ (orange dashed-dotted), $\gamma=0.3$ (magenta dashed-dotted-dotted), $\gamma=0.4$ (cyan short-dashed), and $\gamma=0.5$ (blue short dashed-dotted), in units where $M^{2}_{p}=1$.
-In all cases we have imposed $\Omega_{DE}(x=-\ln(1+z)=0)\equiv\Omega_{DE0}\approx 0.7$ at present, in agreement with observations.
 Figure 4: The evolution of the dark-energy equation-of-state parameter for Barrow holographic dark energy in the Hyperbolic Tangent case with $\Delta(z)=\alpha+\beta\tanh({\gamma z})$, for $\alpha=\beta=\frac{1}{2}$.
@@
 These features reveals the capabilities of the model, and the advantages of Barrow holographic dark energy with running exponent comparing to standard Barrow holographic dark energy.
+We close this subsection by confronting the model of the hyperbolic tangent case at hand with Supernovae type Ia (SNIa) and Cosmic Chronometer data.
+For SNIa, it is well established that the relation between the apparent and absolute magnitudes follows
+$2.5\log\left[\frac{L}{l(z)}\right]=\mu\equiv m(z)-M=5\log\left[\frac{d_{L}(z)_{\text{obs}}}{Mpc}\right]+25~{},$ (4.10)
+with $l(z)$ and $m(z)$ the apparent luminosity and apparent magnitude, and $L$ and $M$ the absolute luminosity and magnitude, respectively, while $d_{L}(z)_{\text{obs}}$ is the luminosity distance.
+Moreover, the theoretical expression for the luminosity distance is given by
+$d_{L}\left(z\right)_{\text{th}}\equiv\left(1+z\right)\int^{z}_{0}\frac{dz^{\prime}}{H\left(z^{\prime}\right)}~{}.$ (4.11)
+Figure 5: The apparent minus absolute magnitude predicted theoretically in the Hyperbolic Tangent case (red - dashed) with $\alpha=\beta=1/2$ and with $\gamma=0.001$, in units where $M_{p}^{2}=1$, on top of the Pantheon SNIa data points from [96].
+For comparison we depict the $\Lambda$CDM curve (black - solid) too.
+Figure 6: The Hubble parameter $H(z)$ in units of Km/s/Mpc as a function of the redshift, for the hyperbolic tangent case with $\alpha=\beta=1/2$ and with $\gamma=0.001$ (red-solid), in $M_{p}^{2}=1$ units, on top of the Cosmic Chronometers data points from [97] at $2\sigma$ confidence level.
+For comparison we depict the $\Lambda$CDM curve (black - solid) too.
+We have imposed $\Omega_{m0}\approx 0.31$.
+Since the hyperbolic tangent case, as well as the $\Lambda$CDM scenario, provide predictions for $H(z)$, we compare the theoretically computed apparent minus absolute magnitudes with the binned Pantheon SNIa data from [96] in Fig.
+5.
+As illustrated, the agreement is very good, with our model accurately reproducing the observed cosmological behavior.
+In addition, the Cosmic Chronometer (CC) dataset, based on the measurement of $H(z)$ from the relative ages of passively evolving galaxies, allows us to compare the observed $H(z)$ values with our model predictions.
+In Fig.
+6, we show the comparison between the theoretical evolution of $H(z)$ in our model with a hyperbolic tangent form and the $\Lambda$CDM model with the CC data points from [97], presented at a $2\sigma$ confidence level.
+The agreement is again very good, with our model reproducing the observed accelerating expansion for the parameter set $\{\alpha,\beta,\gamma\}=\{0.5,0.5,0.001\}$.
+In summary, there exist regions in the parameter space of our model that are able to reproduce the observed evolution of the Hubble function.
+This suggests that this model is a viable candidate.
+A more thorough evaluation, including likelihood analysis and model selection using complete cosmological datasets, will be presented in future work.
 4.5 Trigonometric functions case
@@
 One such case could be the $\Delta(z)=\alpha+\beta\sin{z}$ with $0<\beta<\alpha$, which leads to
-$\!\!\!\!\!\!\!\!\!\!\!\!\!\frac{-(1+z)\Omega_{DE}^{\prime}}{\Omega_{DE}\left(1-\Omega_{DE}\right)}=\sqrt{\Omega_{DE}}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta\sin{z}}{2(\alpha+\beta\sin{z}-2)}}(2-\alpha-\beta\sin{z})$
-$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \,\ +\log{\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]}^{\frac{-(1+z)\beta\cos{z}}{\alpha+\beta\sin{z}-2}}+\alpha+\beta\sin{z}+1,\hskip 32.00002pt$ (4.10)
-$\!\!\!\!\!\!\!\!w_{DE}=-\frac{\alpha+\beta\sin{z}+1}{3}+\frac{(\alpha+\beta\sin{z}-2)\sqrt{\Omega_{DE}}}{3}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta\sin{z}}{2(\alpha+\beta\sin{z}-2)}}$
-$\ \ \ \ \ \ +\frac{(1+z)\beta\cos{z}}{3}\log{\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{1}{2-\alpha-\beta\sin{z}}}}.$ (4.11)
+$\!\!\!\!\!\!\!\!\!\!\!\!\!\frac{-(1+z)\Omega_{DE}^{\prime}}{\Omega_{DE}\left(1-\Omega_{DE}\right)}=\sqrt{\Omega_{DE}}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta\sin{z}}{2(\alpha+\beta\sin{z}-2)}}(2-\alpha-\beta\sin{z})$
+$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \,\ +\log{\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]}^{\frac{-(1+z)\beta\cos{z}}{\alpha+\beta\sin{z}-2}}+\alpha+\beta\sin{z}+1,\hskip 32.00002pt$ (4.12)
+$\!\!\!\!\!\!\!\!w_{DE}=-\frac{\alpha+\beta\sin{z}+1}{3}+\frac{(\alpha+\beta\sin{z}-2)\sqrt{\Omega_{DE}}}{3}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta\sin{z}}{2(\alpha+\beta\sin{z}-2)}}$
+$\ \ \ \ \ \ +\frac{(1+z)\beta\cos{z}}{3}\log{\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{1}{2-\alpha-\beta\sin{z}}}}.$ (4.13)
 Alternatively, one could consider $\Delta(z)=\alpha+\beta\cos{z}$, obtaining
-$\!\!\!\!\!\!\!\!\!\!\!\!\!\frac{-(1+z)\Omega_{DE}^{\prime}}{\Omega_{DE}\left(1-\Omega_{DE}\right)}=\sqrt{\Omega_{DE}}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta\cos{z}}{2(\alpha+\beta\cos{z}-2)}}(2-\alpha-\beta\cos{z})$
-$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\log{\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]}^{\frac{(1+z)\beta\sin{z}}{\alpha+\beta\cos{z}-2}}+\alpha+\beta\cos{z}+1,\hskip 28.00006pt$ (4.12)
-$\!\!\!\!\!\!\!w_{DE}=-\frac{\alpha+\beta\cos{z}+1}{3}+\frac{(\alpha+\beta\cos{z}-2)\sqrt{\Omega_{DE}}}{3}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta\cos{z}}{2(\alpha+\beta\cos{z}-2)}}$
-$\ \ \ \ \ \ \,-\frac{(1+z)\beta\sin{z}}{3}\log{\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{1}{2-\alpha-\beta\cos{z}}}}$ (4.13)
+$\!\!\!\!\!\!\!\!\!\!\!\!\!\frac{-(1+z)\Omega_{DE}^{\prime}}{\Omega_{DE}\left(1-\Omega_{DE}\right)}=\sqrt{\Omega_{DE}}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta\cos{z}}{2(\alpha+\beta\cos{z}-2)}}(2-\alpha-\beta\cos{z})$
+$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\log{\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]}^{\frac{(1+z)\beta\sin{z}}{\alpha+\beta\cos{z}-2}}+\alpha+\beta\cos{z}+1,\hskip 28.00006pt$ (4.14)
+$\!\!\!\!\!\!\!w_{DE}=-\frac{\alpha+\beta\cos{z}+1}{3}+\frac{(\alpha+\beta\cos{z}-2)\sqrt{\Omega_{DE}}}{3}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta\cos{z}}{2(\alpha+\beta\cos{z}-2)}}$
+$\ \ \ \ \ \ \,-\frac{(1+z)\beta\sin{z}}{3}\log{\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{1}{2-\alpha-\beta\cos{z}}}}$ (4.15)
 The behavior of dark energy and matter density parameters, of the deceleration parameter, and of the dark-energy equation-of-state parameter, is similar to the one of the previous models.
@@
 We showed that for increasing $\gamma$ the value of $w_{DE}$ decreases, while for some cases it can experience the phantom-divide crossing, and that in the future $w_{DE}$ can either tend to $-1$ or start increasing again.
-Lastly, for completeness, we also presented the case where the running Barrow exponent has a trigonometric form.
+For completeness, we performed a basic confrontation of the hyperbolic tangent model against the Supernovae type Ia (SnIa) and Cosmic Chronometers (CC) datasets, as a first evidence that they are viable and consistent with observations.
+Lastly, we also presented the case where the running Barrow exponent has a trigonometric form.
 All these features reveal that Barrow holographic dark energy with varying exponent is not only theoretically more justified than the standard, constant-exponent case, but it leads to richer cosmological behavior too.
-It would be interesting to perform observational confrontation with data from Supernovae type Ia (SNIa), Baryonic Acoustic Oscillations (BAO), Cosmic Microwave Background (CMB), and Cosmic Chronometers (CC) probes, in order to extract constraints on the involved parameters, and apply information criteria in order to compare the scenario with the concordance $\Lambda$CDM cosmology.
+It would be interesting to perform a full observational confrontation with data from Supernovae type Ia (SNIa), Baryonic Acoustic Oscillations (BAO), Cosmic Microwave Background (CMB), and Cosmic Chronometers (CC) probes, in order to construct the likelihood contours and extract constraints on the involved parameters, and apply information criteria in order to compare the scenario with the concordance $\Lambda$CDM cosmology.
 Such a full observational analysis is left for a future project.